can you please help me solve this: (2y^3z^2)(6y^4)
Answer by checkley75(3666)

(2Y^3Z^2)(6Y^4) MULTIPLY THE NUMBERS 2*6=12
NOW MULTIPLY THE LIKE TERMS Y^3*Y^4=Y^7.
THUS WE HAVE 12(Y^7)(Z^2) FOR AN ANSWER.